OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle? Answer 23 1 . 3 points

OpenStudy (anonymous):

i think is 265 squarrooted

OpenStudy (turingtest):

it is not, could you show some work?

OpenStudy (anonymous):

11*11 + 12*12

OpenStudy (turingtest):

but the formula is\[a^+b^2=c^2\]where a and b are legs, and c is the hypotenuse, so we need to do something different. We need to solve for one of the legs...

OpenStudy (turingtest):

oops, I meant\[a^2+b^2=c^2\]

OpenStudy (anonymous):

so we use the a-c

OpenStudy (anonymous):

so 23

OpenStudy (diyadiya):

C^2-a^2

OpenStudy (anonymous):

squared

OpenStudy (anonymous):

23

OpenStudy (diyadiya):

12^2-11^2 144-121=23 Yes Right

OpenStudy (anonymous):

it is squared or isnt

OpenStudy (turingtest):

don't forget to take the square root. yes

OpenStudy (anonymous):

ok thank lots guys

OpenStudy (diyadiya):

Oh yes \[\sqrt{23}\]